OpenStudy (dj3strella):

Can someone help me please I forgot how to do this problem 12. If m varies directly with n and m=6 when n=9, what is the value of n when m=8 a.) n=24 b.) n=12 c.) n=16/3 d.) n=24/4

OpenStudy (dj3strella):

I don't think theres a typo... @Hero

OpenStudy (dj3strella):

yes @hero i found a typo

OpenStudy (dj3strella):

@dan815 can you help me?

OpenStudy (dan815):

km=n 9/6 * 6 = 9 when m=8 n=... 9/6 * 8 = 24/2 = 12

OpenStudy (dj3strella):

so it equals 12?